OpenStudy (sarahvubble):

How to do this ? I got help for 1&2 but I don't know how this works . I've been sick from school so I just don't get how this is done ? Please help me . Thank you

OpenStudy (shadowlegendx):

When you get back...what is your question?

OpenStudy (sarahvubble):

How to add and subtract w like denominators by using LCD??

OpenStudy (shadowlegendx):

Like \[\frac{ 1 }{ 2 } - \frac{ 1 }{ 4 } \]

OpenStudy (shadowlegendx):

?

OpenStudy (shadowlegendx):

The way you would solve the above problem is by finding the Lowest Common Denominator, or LCD. 2 goes into 4 twice, 4 goes into 4 once, 4 can be our LCD. We want to convert 1/2 to a number over 4. Since we multiply 2 on the bottom, by 2, to get 4, we do the same to 1 on the top. \[\frac{ 2 }{ 2 } \times \frac{ 1 }{ 2 } = \frac{ 2 }{ 4 }\]

OpenStudy (shadowlegendx):

Now that we have our common denominator. \[\frac{ 2 }{ 4 } - \frac{ 1 }{ 4 } = \frac{ 1 }{ 4 }\]

OpenStudy (shadowlegendx):

Remember that only the numerators subtract(the number on the top), not the denominators

OpenStudy (turbostudier):